The potential at the surface of a sphere (radius R) is given by

V_{0}=k \cos 6 \theta

V0

​=kcos6θ ,

where k is a constant. Find the potential inside and outside the sphere, as well as the surface charge density σ (θ) on the sphere. (Assume there’s no charge inside or outside the sphere.)


1
Expert's answer
2022-02-17T11:45:21-0500

We know that

Poisson equation

"\nabla^2V=-\frac{\rho}{\epsilon_0}"

Charge density is zero r>R,r<R

"\rho=0"

"\nabla^2V=0"

"V(r,\theta)=\Sigma A_l r^lP_l(cos\theta)" r<R

"V(r,\theta)=\Sigma B_l r^{-l+1}P_l(cos\theta)" r>R

"V_0=kcos6\theta"

Comparison the gives potential with legendre's polynomial equation

"V(r,\theta)=kP_l(cos6\theta)"

"A_1=\frac{k}{R}\\B_1={kR^2}"

Now

"V_1(r_1,\theta)=\frac{kR^2}{r^2}cos6\theta" r>R

"V(r,\theta)=\frac{Kr}{R}cos6\theta" r<R